a cat, meerkat and rat are fighting over a bat. the cat is pulling with a force of 25N (south), the rat is pulling with a force of 12N (North 20 degrees east) and meerkat is pulling with a force of 18N (N30 degrees west) what is the net force and acceleration acting on the 2.2kg bat


1
Expert's answer
2021-01-19T20:00:36-0500

"F_x=-18\cdot\cos30°+12\cdot \cos20°=-4.31(N)"


"F_y=18\cdot\sin30°+12\cdot \sin20°-25=-11.89(N)"


"F=\sqrt{11.89^2+4.31^2}=12.65(N)" . Answer


"a=F\/m=12.65\/2.2=5.75(m\/s^2)" . Answer